### v1

# Axion Signal Search Using Hybrid Nuclear–Electronic Spin Systems

Abstract
Conventional nuclear magnetic resonance searches for the galactic axion wind lose sensitivity at low frequencies due to the unfavourable scaling of inductive readout. Here, we propose a hybrid architecture where the hyperfine interaction transduces axion-driven nuclear precession into a high-bandwidth electron-spin readout channel. We demonstrate analytically that this dispersive upconversion preserves the specific sidereal and annual modulation signatures required to distinguish dark matter signals from instrumental backgrounds. When instantiated in a silicon $^{209}\text{Bi}$ donor platform, the hybrid sensor is projected to outperform direct nuclear detection by more than an order of magnitude over the $10^{-16}-10^{-6}\text{eV}$ wide mass range. With collective enhancement, the design reaches a $5\sigma$ sensitivity to DFSZ axion-nucleon couplings within one year, establishing hyperfine-mediated sensing as a competitive path for compact, solid-state dark matter searches.
Introduction- Searches for ultralight dark matter such as axions and axionlike particles (ALPs) have attracted increasing attention from the quantum sensing and particle physics communities [34, 5, 24, 37, 31, 15, 10, 18, 44, 1, 2, 3, 20, 36, 6, 23, 7, 32, 8, 59]. For axions and ALPs with derivative couplings to nucleon spins, the interaction can be expressed as an effective axion wind that coherently drives nuclear moments at a narrowband Compton frequency $\omega_{a}\simeq m_{a}c^{2}/\hbar$, with sidereal and annual modulations set by the laboratory’s motion through the Galactic frame [25, 49]. These direction dependent kinematic signatures provide a handle for discriminating axion signals from terrestrial backgrounds.
Current nuclear magnetic resonance (NMR) experiments, exemplified by the CASPEr-Wind program, have used large nuclear polarizations and long spin coherence times to set leading bounds on such axion–nucleon couplings [9, 19, 57]. However, conventional inductive NMR readout faces a fundamental limitation at low axion Compton frequencies. Because the pickup voltage scales as $V\propto\omega_{a}$ and the nuclear gyromagnetic ratio $\gamma_{N}$ is small, the sensitivity degrades rapidly for lighter axions, even before accounting for bandwidth, resonator quality factor, or magnetometer noise. This unfavourable low-frequency scaling makes the sub-Hz to kHz axion mass window particularly challenging for NMR-based wind searches.
Semiconductor spin qubits provide a complementary route. Compared with nuclei, electron spins have a much larger gyromagnetic ratio and benefit from mature, high-sensitivity readout, including spin-to-charge conversion via Pauli spin blockade and gate- or microwave-based dispersive reflectometry [16, 26, 14, 12]. In donor and quantum dot platforms the strong hyperfine interaction $A\bm{I}\cdot\bm{S}$ tightly couples nuclear and electronic degrees of freedom, enabling nuclear dynamics to be mapped onto charge or electron-spin observables with high fidelity [40, 42, 48, 47]. This motivates hybrid sensors in which a nuclear spin ensemble serves as an axion-sensitive antenna, while a coupled electron spin provides a fast, high-bandwidth readout channel. The central challenge is to bridge the disparate nuclear and electronic timescales and energy scales, and to transduce the nuclear response into the electronic domain without degrading the spatiotemporal coherence or sidereal and annual signatures.
In this manuscript, we address this challenge by developing a hyperfine-mediated upconversion readout for axion wind searches. Building on Ref. [51, 52] on single electron and quantum dot based axion sensing, we extend this method to a nuclear–electronic hybrid architecture with collective nuclear enhancement. First, we construct a filter function formalism for nuclear–electronic transduction, defining a hybrid gain $G_{\text{hyb}}$ and nuclear filter function $Y_{N}(\omega)$ that give platform-independent expressions for sensitivity. Second, we show that hyperfine upconversion preserves the axion wind kinematics: after demodulation, the electronic readout exhibits a sidereal triplet centered at the sidereal frequency $\Omega_{\star}$ with annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$, providing a geometric veto against terrestrial systematics. Finally, we validate the approach in a silicon $^{209}\text{Bi}$ donor system, incorporating realistic coherence times, control pulse constraints, and readout noise to benchmark the performance against state-of-the-art NMR strategies.
We find that the hybrid readout achieves order-of-magnitude improved sensitivity over direct nuclear detection across the accessible mass range. With collective nuclear-spin enhancement and a high-$Q$ resonator, the architecture reaches $5\sigma$ sensitivity to DFSZ-scale axion-nucleon coupling $g_{aNN}$ within one year of integration. Previous axion wind detection schemes suffer from fundamental readout-scaling limits especially at low masses. We show that hyperfine-mediated nuclear-to-electron transduction eliminates this $\omega_{a}$ dependence, enabling sensitivity in a previously inaccessible $10^{-16}–10^{-6}$ eV window.
Model and theory We consider a semiconductor donor system with a spin-1/2 electron $\bm{S}$ and a nuclear spin $\bm{I}$ (for 209Bi, $I=9/2$) in a static magnetic field $\bm{B}=B_{z}\hat{\text{z}}$ [56, 58]. Setting $\hbar=1$, the system Hamiltonian in the laboratory frame is given by:
$H_{\text{total}}=\omega_{e}S_{z}+\omega_{N}I_{z}+A\bm{S}\cdot\bm{I}+H_{aN}(t).$ (1)
The electron and nuclear Larmor frequencies are $\omega_{e}=|\gamma_{e}|B_{z}$ and $\omega_{N}=\gamma_{N}B_{z}$, where $\gamma_{e}$ is the electron gyromagnetic ratio. The isotropic hyperfine interaction constant $A$ is taken from experiment. The last term encodes the derivative coupling of an axion like particle field to nucleons and yields an effective axion wind interaction: $H_{aN}(t)=-\gamma_{N}\bm{I}\cdot\bm{B}_{a}(t)$, where $\bm{B}_{a}(t)$ is the axion induced magnetic field, which can be expressed as $g_{aNN}\nabla a/(2m_{N}\gamma_{N})$. Treating the axion field as a coherent classical wave, we write the effective field as $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos\left(\omega_{a}t+\phi\right)$, where $\omega_{a}\simeq m_{a}$ is the Compton frequency. The spatial gradient $\nabla a$ fixes the orientation of $\bm{B}_{a}$ in the celestial frame, which manifests as sidereal and annual modulations of the signal amplitude in the laboratory frame due to Earth’s motion. The coherence time of this drive is characterized by $\tau_{a}\simeq 2\pi/\left(m_{a}v_{0}^{2}\right)$ under the SHM.
We focus on the longitudinal component of the axion-induced field, $B_{a,z}(t)$, which modulates the nuclear Larmor frequency. To probe axion signals across a broad mass range, we apply the dynamical decoupling sequences that tailor the nuclear spectral response [39, 55, 17]. Specifically, we use Ramsey free induction decay (FID), Hahn echo, XY8, and Carr-Purcell-Meiboom-Gill (CPMG) protocols. The spectral sensitivity of a sequence is characterized by its filter function $Y_{N}(\omega)$, the Fourier transform of the time-domain sensitivity function $y_{N}(t)\in\{0,\pm 1\}$ over the sensing duration. For a narrowband axion field $B_{a,z}(t)=B_{a}\cos(\omega_{a}t+\phi)$, the accumulated phase $\varphi_{N}$ at the end of the sequence is given by:
$\varphi_{N}\approx\gamma_{N}B_{a}|Y_{N}(\omega_{a})|\cos(\omega_{a}t_{0}+\Phi_{\text{off}})\,$, (2)
where $t_{0}$ marks the start of the sequence and $\Phi_{\text{off}}$ includes the phase of the filter function and the axion field. The choice of sequence determines the detection window. The Ramsey ($y_{N}=1$) exhibits a low-pass filter response $|Y_{N}(\omega)|\approx\tau\operatorname{sinc}(\omega\tau/2)$, maximizing sensitivity to ultralight axion fields. Conversely, Hahn echo and CPMG sequences act as band-pass filters, suppressing low-frequency noise while extending sensitivity to higher-frequency (oscillating) axion fields determined by the pulse spacing [13, 4, 43].
Experimentally, the nuclear spin is initialized in a superposition state (e.g., along $\hat{\text{x}}$). After accumulating the axion-induced phase $\varphi_{N}$ over time $\tau$, a final $\pi/2$ pulse maps this transverse phase onto the longitudinal polarization $\langle I_{z}\rangle$. In the small-signal limit ($|\varphi_{N}|\ll 1$), the observable signal is proportional to the phase, $\delta\langle I_{z}\rangle\approx I\varphi_{N}$, allowing direct readout of the axion interaction. In the high-field limit, where $|\omega_{e}-\omega_{N}|\gg A$, we apply the secular approximation to neglect the terms $A(S_{+}I_{-}+S_{-}I_{+})/2$, which only leads to negligible energy shifts (See supplementary material).
The effective Hamiltonian $H_{\text{eff }}=(\omega_{e}+AI_{z})S_{z}+\omega_{N}I_{z}$ indicates that the nuclear polarization acts as a local field modulating the electron Larmor frequency. Consequently, the axion-driven nuclear dynamics are dispersively upconverted into an electron frequency modulation (FM): $\delta\omega_{e}(t)=A_{\text{eff}}|\gamma_{N}Y_{N}(\omega_{a})B_{a,z}|\cos(\omega_{a}t+\phi)$. Here, the signal amplitude is proportional to the nuclear filter gain $|Y_{N}|$, and we introduce $A_{\text{eff}}=AG_{\text{drv}}$ to generalize the static hyperfine coupling to driven schemes like Hartmann-Hahn resonance or Floquet dressing, where the effective interaction strength is dynamically enhanced.
It is convenient to express the axion induced electron FM as an effective magnetic field in the electron channel, $B_{\text{eff}}^{(e)}\equiv\delta\omega_{e,0}/\gamma_{e}$. The full axion-nucleus-electron transduction chain is summarized into a hybrid gain, defined by $B_{\text{eff}}^{(e)}/|B_{a,z}|$:
$G_{\text{hyb}}(\omega_{a})=\frac{AI}{\gamma_{e}}\gamma_{N}|Y_{N}(\omega_{a})|G_{\text{drv}}\left[1+\mathcal{O}\left(\frac{A^{2}}{\Delta^{2}}\right)\right]\,$, (3)
where $G_{\text{drv}}=1$ in the absence of dressing or driving. The electron signal is enhanced by long nuclear coherence (large $|Y_{N}|$) and strong hyperfine coupling $A$. In an electron Ramsey or echo probe with sensitivity function $y_{e}(t)$ and sinusoidal frequency modulation $\delta\omega_{e}(t)=\delta\omega_{e,0}\cos(\omega_{a}t)$, continuous readout yields Bessel sidebands at $\omega_{e}\pm n\omega_{a}$ with modulation index $\beta=\delta\omega_{e,0}/\omega_{a}$. The axion wind direction $\hat{\bm{u}}_{a}$ is fixed in the Galactic frame while the laboratory quantization axis rotates with Earth, so the projection $R(t)\hat{\bm{u}}_{a}\cdot\hat{\bm{z}}$ generates a sidereal component at $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$ [52]. After upconversion, the modulation spectrum retains the characteristic sidereal triplet and annual sidebands, providing an instrument-independent astrophysical signature.
Having established how the axion signal maps onto an effective electron field and its modulation pattern, we now quantify the resulting signal-to-noise performance in the electron and nuclear channels. We characterize the SNR in the electron and nuclear channels by three time scales: the electron/nuclear coherence time $T_{2}^{(e,N)}$, the axion coherence time $\tau_{a}$ set by the Galactic velocity dispersion, and the total observation time $T_{\text{obs}}$. The effective coherent integration time $T_{\text{coh}}^{(e,N)}=\min\{T_{2}^{(e,N)},\tau_{a},T_{\text{obs}}\}$ then sets how the SNR in each channel scales with the effective signal field and the corresponding readout noise. In the Ramsey limit $\omega_{a}\tau_{N}\ll 1$, the nuclear response is approximately flat over the interrogation time, while echo/CPMG sequences produce narrowband peaks. Long nuclear coherence, strong hyperfine coupling, and low electron readout noise allow the upconversion scheme to surpass direct nuclear readout whenever $T_{2}^{(e)}$ and $\tau_{a}$ are large enough to support coherent integration. Including both readout noise floors, the relative sensitivity is
$\frac{\text{SNR}_{e}}{\text{SNR}_{N}}=G_{\text{hyb}}\frac{\eta_{B}^{(N)}}{\eta_{B}^{(e)}}\sqrt{\frac{T_{\text{coh}}^{(e)}}{T_{\text{coh}}^{(N)}}}\,.$ (4)
Thus hybrid electron readout with matched filtering exceeds direct nuclear spin detection when the hyperfine gain is appreciable, the electron readout noise is lower, and the electron coherence time is comparable to the nuclear coherence time.
Figure 1: Schematic of a prototypical experimental design. A gate-defined electron spin qubit (left) provides a fast, high-bandwidth readout channel, while a donor nuclear spin (right) serves as the axion-sensitive element. Gates P1, P2, Q1, and Q2 define the confinement potentials and are shown for illustration. Realizations can follow electron- and hole-based architectures in Refs. [46, 41, 30, 28, 35, 27, 29], with the donor nuclear spin acting as the probe of the axion wind field as in Refs. [22, 53, 54, 38, 11, 50].
Experimental prototype The device shown in Fig. 1 is a chip-integrated hybrid nuclear-electronic spin sensor designed for axion/ALPs search. All functional elements-donor ensemble, control gates, and readout structures are integrated on a single chip. The chip is enclosed in radio-frequency and magnetic shielding to suppress stray fields and is thermally anchored to mK stage that provide stability for initialization, manipulation, and readout [21, 33]. Superconducting NbTi traces and twisted-pair wiring are used within the EM shielded enclosure to minimize loss, reduce Johnson noise/AC noise, and maintain phase coherence across the hybrid nuclear-electronic network.
Results To estimate the axion-search sensitivity of the proposed sensors, we normalize the hybrid gain of each isotope to the sensitivity baseline introduced in Ref. [51]. Building on Eq. (3), we evaluate the electron phase filter $F_{e}(\omega)$ and the nuclear filter $Y_{N}(\omega)$ for a Ramsey/Hahn/CPMG sequence bank and determine the corresponding scan bands from their passbands. As expected from Eq. 4, the strong hyperfine coupling in 209Bi produces the largest hybrid gain at fixed $\tau_{N}$ and corresponding sensitivity advantage.
The hybrid spectral response can be written as $H(\omega;\tau,N)=F_{e}(\omega;\tau)\,Y_{N}(\omega;\tau,N)$, where $\tau$ is the sensing duration and $N$ the number of $\pi$ pulses. The electron filter $F_{e}$ suppresses low-frequency components and imposes fixed spectral zeros [Fig. 2(a)], while the nuclear filter $Y_{N}$ sets the tunable passband. Under CPMG control, the main passband is centered near $f_{c}\simeq N/\tau$ with width of order $1/\tau$ and quality factor $Q\sim N$ [Fig. 2(b)], enabling narrowband frequency targeting without sacrificing coherence. Together, $F_{e}$ and $Y_{N}$ define a controllable detection band that supports systematic scanning of the axion mass range.
Figure 2: (a) Normalized electron filter functions $|F(\omega)|^{2}/\tau^{2}$ for Ramsey and Hahn echo versus normalized frequency $\xi=f\tau$; the Hahn sequence narrows the main lobe and shifts the first zero from $\xi\simeq 1$ (Ramsey) to $\xi\simeq 2$. (b) Nuclear toggling filter amplitude $|Y_{N}(\omega)|/\tau$ for Ramsey, Hahn, and CPMG with $N=4,8,16$, exhibiting passbands centered near $\xi\simeq N$ with width $\sim 1/\tau$. This filter structure enables tunable frequency selectivity while preserving the sidereal modulation pattern, in contrast to conventional inductive NMR. Together these properties define the accessible mass-scan range.
Within this bounded window, matched filtering yields a resolved hybrid response with a signal-to-noise ratio scaling as $\mathrm{SNR}\propto A\sqrt{\tau}|H(\omega_{a})|$, where $H(\omega)$ denotes the joint electron–nuclear spectral filter. This structure enables mass scanning without degrading the sidereal modulation fingerprint of the axion wind. A full derivation of the scan constraints, pulse-selection rules, and the explicit bandwidth expressions used to generate Fig. 2 is provided in the Supplementary Material.
This scaling explains the monotonic trends also in Supplementary Material: Fig. 1, where longer $T_{2,N}$ directly lifts the hybrid gain, while larger hyperfine $A$ (e.g., 209Bi) raises the transduction into the electron readout channel. Consistent with the modulation structure discussed earlier, the joint filter $H(\omega_{a})$ response preserves the same spectral fingerprint and therefore remains compatible with matched-filter detection.
Figure 3: (a) Time-domain evolution of the electron rotating-frame longitudinal magnetization $\langle S_{x}(\tau)\rangle$ under continuous spin lock at a Rabi frequency $\Omega_{R}/2\pi=50~\text{kHz}$. The axion-induced nuclear precession introduces a weak frequency modulation at $\omega_{a}/2\pi=5~\text{kHz}$. (b) Power spectral density of the same signal, $|\tilde{S}_{x}(\omega)|^{2}$, showing the locked carrier at $\Omega_{R}/2\pi$ (black dashed line) and symmetric axion-induced sidebands at $\Omega_{R}/2\pi\pm\omega_{a}/2\pi$ (red dashed lines). These spectral sidebands represent the frequency-encoded axion response characteristic of spin-lock detection.
While the pulsed protocols discussed above rely on accumulated phase to extract the axion-induced nuclear response, the same dynamics can also be accessed through a continuous-drive scheme. In this regime, the electronic qubit is held in a spin-locked state, so the axion-driven nuclear precession appears as a frequency modulation of the driven electron rather than a freely evolving phase. In the noiseless illustrative signal shown in Fig. 3, this manifests as weak modulation in the rotating-frame trajectory and as symmetric sidebands around the Rabi carrier, demonstrating the frequency-encoded detection mechanism without requiring a final phase-mapping pulse. This readout is not only a qualitative signature but also determines the architecture’s achievable sensitivity, since the spin-lock mode remains compatible with coherent integration and matched-filter analysis. The projected sensitivity enabled by this readout strategy is shown in Fig. 4.
Figure 4: Projected $5\sigma$ sensitivity to axion-nucleon couplings. Shown are the calculated discovery thresholds on $g_{aNN}$ for a 209Bi donor ensemble under several standard magnetometry protocols: Ramsey (Blue), Hahn-echo (Orange), XY8 (Green), and continuous spin-lock ($T_{1\rho}$) detection (Black). All curves include realistic device parameters and noise models (Supplementary Tables II–III), and assume collective enhancement from an entangled ensemble of $N=10^{6}$ nuclear spins together with resonator gain $G_{\text{res}}(Q{=}10^{5})$. The effective signal field entering each protocol is $B_{e,\text{eff}}=F_{\text{ent}}G_{\text{res}}G_{\text{hyb}}B_{a}^{(N)}$. The purple shaded region represents the spin-lock Monte-Carlo distribution (median and 5%-95% band), which provides the best sensitivity across most of the mass range. The yellow band denotes the DFSZ model predictions, and the horizontal magenta region indicates the current SN1987A cooling bounds (95% C.L.). Part of the parameter windows are created from Ref. [45].
As determined by the filter-architecture constraints in Fig. 2, the achievable search region for the proposed hybrid platform is shown in Fig. 4. The Coloured band marks the portion of parameter space that is simultaneously compatible with the spectral passband and the hardware operating limits, representing the practically scannable mass window of the architecture. This window spans axion Compton frequencies from the sub-Hz regime to the kHz range, a range where conventional NMR-based wind searches are limited by inductive readout scaling. Regions outside this band are not eliminated physically, but lie outside the spectral response or operational range of the present implementation. The emergence of a contiguous, tunable, and readout-compatible scan band highlights a key advantage of the hybrid approach: it enables efficient access to the ultralight regime while maintaining signal contrast and spectral resolution, unlike purely nuclear or purely electronic sensing strategies.
Conclusion Hybrid nuclear-electronic upconversion enables axion wind sensitivity beyond the nuclear-inductive limit while preserving the required directional signatures. This approach is compatible with existing solid-state qubit platforms, enabling experimental validation without new fundamental infrastructure. By combining an axion-coupled nuclear ensemble with electronic readout, the architecture exploits strong axion-nucleus interactions, long nuclear coherence, and hyperfine-mediated signal transfer to a fast, high-fidelity electron channel. The framework links coherence times, hyperfine strength, and technical noise to achievable performance and identifies the sidereal modulation pattern as a robust detection fingerprint. Simulations based on silicon donor platforms show that bismuth, with its large hyperfine coupling, yields the greatest hybrid gain, and with collective enhancement and resonator coupling the scheme reaches projected $5\sigma$ DFSZ-level sensitivity across $m_{a}=10^{-16}$-$10^{-6}\,\text{eV}$ with Monte Carlo simulation. The approach is compatible with CMOS fabrication, scalable donor arrays, and parallel readout, and supports projected sensitivity down to $g_{aNN}\sim 10^{-25}$, positioning this hybrid solid-state platform as a competitive route toward next-generation dark-matter searches.