### v2

# Axion Signal Search Using Hybrid Nuclear–Electronic Spin Systems

Abstract
Conventional nuclear magnetic resonance searches for the galactic axion wind face technical challenges at lower frequencies, where the sensitivity of inductive readout scales less efficiently. Here, we propose a hybrid architecture where the hyperfine interaction transduces axion-driven nuclear precession into a high-bandwidth electron-spin readout channel. We demonstrate analytically that this dispersive upconversion preserves the specific sidereal and annual modulation signatures required to distinguish dark matter signals from instrumental backgrounds. When instantiated in a silicon 209Bi donor platform, the hybrid sensor is projected to outperform direct nuclear detection by more than an order of magnitude over the $10^{-16}-10^{-6}$ eV wide mass range. With additional collective and resonant enhancements, the design can reach discovery-level sensitivity to DFSZ-scale axion–nucleon couplings with integration over measurement time, establishing hyperfine-mediated sensing as a compact solid-state path for dark-matter searches.
Introduction- Searches for ultralight dark matter, including axions and axion-like particles (ALPs), have attracted broad interest in quantum sensing and particle physics [35, 6, 25, 38, 32, 16, 11, 19, 45, 2, 3, 4, 21, 37, 7, 24, 8, 33, 9, 60]. For axions/ALPs with derivative couplings to nucleon spins, the field acts as an effective axion wind that drives nuclear spins at the Compton frequency $\omega_{a}\simeq m_{a}c^{2}/\hbar$ and produces sidereal and annual modulations in the laboratory frame [26, 50]. These directional modulations provide an astrophysical test ground against terrestrial backgrounds.
NMR searches, exemplified by CASPEr-Wind, exploit large nuclear polarization and long coherence to set leading bounds on axion–nucleon wind couplings [10, 20, 58]. Axion-driven sample magnetization can be read out inductively, where the pickup voltage scales as $V\propto\mathrm{d}\Phi/\mathrm{d}t\propto\omega_{a}$, or via precision magnetometry (e.g., SQUIDs or atomic magnetometers) that measures flux or field more directly. In the coherent averaging regime, the optimal coherent integration time is limited by the axion coherence time $\tau_{a}\propto 1/\omega_{a}$, so the explicit $\omega_{a}$ in inductive voltage pickup partly cancels and, for an otherwise white noise floor, $\mathrm{SNR}\propto\omega_{a}\sqrt{\tau_{a}}\sim\sqrt{\omega_{a}}$. Compared with nuclei, electron spins have a much larger gyromagnetic ratio and benefit from mature, high-sensitivity readout, including spin-to-charge conversion via Pauli spin blockade and gate or microwave-based dispersive reflectometry [17, 27, 15, 13]. In donor and quantum dot platforms the strong hyperfine interaction $A\bm{I}\cdot\bm{S}$ tightly couples nuclear and electronic degrees of freedom, enabling nuclear dynamics to be mapped onto charge or electron-spin observables with high fidelity [41, 43, 49, 48]. This motivates hybrid sensors in which a nuclear ensemble acts as the axion-sensitive element while a coupled electron provides fast readout. The central challenge is to bridge the disparate nuclear and electronic timescales and energy scales, and to transduce the nuclear response into the electronic domain without degrading the spatiotemporal coherence or sidereal and annual signatures.
Here we develop a hyperfine-mediated upconversion readout for axion-wind searches. As a benchmark, we quote sensitivity relative to the DFSZ (Dine–Fischler–Srednicki–Zhitnitsky) class of QCD axion models, which arise in two-Higgs-doublet extensions of the Standard Model and predict axion couplings to Standard-Model fermions set by the Peccei–Quinn scale. In this work we target the axion–nucleon wind coupling that drives nuclear spins; the electron spin is used as a transduction/readout channel via hyperfine-mediated upconversion rather than as the primary axion-coupled sensor. Building on Refs. [52, 53], we extend the approach to a nuclear–electronic hybrid architecture with collective enhancement. First, we construct a filter function formalism for nuclear–electronic transduction, defining a hybrid gain $G_{\text{hyb}}$ and nuclear filter function $Y_{N}(\omega)$ that give platform-independent expressions for sensitivity. Second, we show that hyperfine upconversion preserves the axion wind kinematics: after demodulation, the electronic readout exhibits a sidereal triplet centered at the sidereal frequency $\Omega_{\star}$ with annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$, providing a geometric veto against terrestrial systematics. Finally, we validate the approach in a silicon $^{209}\text{Bi}$ donor system, incorporating realistic coherence times, control pulse constraints, and readout noise to benchmark the performance against state-of-the-art NMR strategies. We highlight Si:209Bi donors because their large isotropic hyperfine coupling enables strong transduction together with established control and long coherence in enriched silicon. This combination makes Si:Bi particularly advantageous for hyperfine-mediated upconversion.
We find that the hybrid readout achieves order-of-magnitude improved sensitivity over direct nuclear detection across the accessible mass range. With collective nuclear-spin enhancement and a high-$Q$ resonator, the architecture reaches $5\sigma$ sensitivity to DFSZ-scale axion-nucleon coupling $g_{aNN}$ within one year of integration. At low masses the overall reach of wind searches is often limited by the readout chain rather than by nuclear coherence alone, particularly for pickup-coil voltage detection with its explicit $\omega_{a}$ transduction factor. Hyperfine-mediated nuclear-to-electron transduction removes this explicit inductive penalty by mapping the axion-driven nuclear response onto a high-bandwidth electronic readout channel, leaving the remaining frequency dependence set by $\tau_{a}$ and by the control-sequence filter functions.
Model and theory- We consider a semiconductor donor system with a spin-1/2 electron $\bm{S}$ and a nuclear spin $\bm{I}$ (for 209Bi, $I=9/2$) in a static magnetic field $\bm{B}=B_{z}\hat{\text{z}}$ [57, 59]. The laboratory frame Hamiltonian is:
$H_{\text{total}}=\omega_{e}S_{z}+\omega_{N}I_{z}+A\bm{S}\cdot\bm{I}+H_{aN}(t).$ (1)
The electron and nuclear Larmor frequencies are $\omega_{e}=|\gamma_{e}|B_{z}$ and $\omega_{N}=\gamma_{N}B_{z}$, where $\gamma_{e}$ is the electron gyromagnetic ratio. The isotropic hyperfine interaction constant $A$ is taken from experiment. The axion–nucleon interaction is written as an effective field $H_{aN}(t)=-\gamma_{N}\bm{I}\cdot\bm{B}_{a}(t)$, with $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos(\omega_{a}t+\phi)$. Here $\bm{B}_{a}(t)$ is the axion induced magnetic field, which can be expressed as $g_{aNN}\nabla a/(2m_{N}\gamma_{N})$. Treating the axion field as a coherent classical wave, we write the effective field as $\bm{B}_{a}(t)=\bm{B}_{a,0}\cos\left(\omega_{a}t+\phi\right)$, where $\omega_{a}\simeq m_{a}$ is the Compton frequency. The fixed Galactic-frame direction of $\bm{B}_{a}$ leads to sidereal and annual modulations in the lab frame, and the drive remains coherent for a time $\tau_{a}$ set by the velocity dispersion. The coherence time of this drive is characterized by $\tau_{a}\simeq 2\pi/\left(m_{a}v_{0}^{2}\right)$ under the SHM.
We focus on the longitudinal component of the axion-induced field, $B_{a,z}(t)$, which modulates the nuclear Larmor frequency. To cover a broad frequency range, we use standard Ramsey/echo and multipulse sequences (e.g., Ramsey, Hahn echo, XY8, and CPMG) to shape the nuclear spectral response [40, 56, 18]. The spectral sensitivity of a sequence is characterized by its filter function $Y_{N}(\omega)$, the Fourier transform of the time-domain sensitivity function $y_{N}(t)\in\{0,\pm 1\}$ over the sensing duration. The values $\pm 1$ corresponds to free precession segments whose sign is toggled by pulses, while 0 accounts for intervals with negligible phase accumulation. For a narrowband axion field $B_{a,z}(t)=B_{a}\cos(\omega_{a}t+\phi)$, the accumulated phase $\varphi_{N}$ at the end of the sequence is given by:
$\varphi_{N}\approx\gamma_{N}B_{a}|Y_{N}(\omega_{a})|\cos(\omega_{a}t_{0}+\Phi_{\text{off}})\,$, (2)
where $t_{0}$ marks the start of the sequence and $\Phi_{\text{off}}$ includes the phase of the filter function and the axion field. The choice of sequence determines the detection window. The Ramsey ($y_{N}=1$) exhibits a low-pass filter response $|Y_{N}(\omega)|\approx\tau\operatorname{sinc}(\omega\tau/2)$, maximizing sensitivity to ultralight axion fields. Conversely, Hahn echo and CPMG sequences act as band-pass filters, suppressing low-frequency noise while extending sensitivity to higher-frequency (oscillating) axion fields determined by the pulse spacing [14, 5, 44].
Experimentally, the nuclear spin is initialized in a superposition state (e.g., along $\hat{\text{x}}$). After accumulating the axion-induced phase $\varphi_{N}$ over time $\tau$, a final $\pi/2$ pulse maps this transverse phase onto the longitudinal polarization $\langle I_{z}\rangle$. In the small-signal limit ($|\varphi_{N}|\ll 1$), the observable signal is proportional to the phase, $\delta\langle I_{z}\rangle\approx I\varphi_{N}$, allowing direct readout of the axion interaction. In the dispersive regime the hyperfine interaction produces a longitudinal shift of the electron splitting proportional to $I_{z}$, enabling nuclear-to-electron transduction [1].
The effective Hamiltonian $H_{\text{eff}}=(\omega_{e}+AI_{z})S_{z}+\omega_{N}I_{z}$ indicates that the nuclear polarization acts as a local field modulating the electron Larmor frequency. Consequently, the axion-driven nuclear dynamics are dispersively upconverted into an electron frequency modulation (FM): $\delta\omega_{e}(t)=A_{\text{eff}}I|\gamma_{N}Y_{N}(\omega_{a})B_{a,z}|\cos(\omega_{a}t+\phi)$. Here, the signal amplitude is proportional to the nuclear filter gain $|Y_{N}|$, and we introduce $A_{\text{eff}}=AG_{\text{drv}}$ to generalize the static hyperfine coupling to driven schemes like Hartmann-Hahn resonance or Floquet dressing, where the effective interaction strength is dynamically enhanced.
We express the transduced signal as an effective electron field $B_{\mathrm{eff}}^{(e)}\equiv\delta\omega_{e,0}/\gamma_{e}$ and define the hybrid gain:
$G_{\text{hyb}}(\omega_{a})=\frac{AI}{\gamma_{e}}\gamma_{N}|Y_{N}(\omega_{a})|G_{\text{drv}}\left[1+\mathcal{O}\left(\frac{A}{2\Delta}\right)\right]\,$, (3)
where $G_{\text{drv}}=1$ in the absence of dressing or driving. The electron signal is enhanced by long nuclear coherence (large $|Y_{N}|$) and strong hyperfine coupling $A$. In an electron Ramsey or echo probe with sensitivity function $y_{e}(t)$ and sinusoidal frequency modulation $\delta\omega_{e}(t)=\delta\omega_{e,0}\cos(\omega_{a}t)$, continuous readout yields Bessel sidebands at $\omega_{e}\pm n\omega_{a}$ with modulation index $\beta=\delta\omega_{e,0}/\omega_{a}$. The axion wind direction $\hat{\bm{u}}_{a}$ is fixed in the Galactic frame while the laboratory quantization axis rotates with Earth, so the projection $R(t)\hat{\bm{u}}_{a}\cdot\hat{\bm{z}}$ generates a sidereal component at $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$ [53, 1]. After upconversion, the modulation spectrum retains the characteristic sidereal triplet and annual sidebands, providing an instrument-independent astrophysical signature.
Having established how the axion signal maps onto an effective electron field and its modulation pattern, we now quantify the resulting signal-to-noise performance in the electron and nuclear channels. We characterize the SNR in the electron and nuclear channels by three time scales: the electron/nuclear coherence time $T_{2}^{(e,N)}$, the axion coherence time $\tau_{a}$ set by the Galactic velocity dispersion, and the total observation time $T_{\text{obs}}$. The effective coherent integration time $T_{\text{coh}}^{(e,N)}=\min\{T_{2}^{(e,N)},\tau_{a},T_{\text{obs}}\}$ then sets how the SNR in each channel scales with the effective signal field and the corresponding readout noise. In the Ramsey limit $\omega_{a}\tau_{N}\ll 1$, the nuclear response is approximately flat over the interrogation time, while echo/CPMG sequences produce narrowband peaks. Long nuclear coherence, strong hyperfine coupling, and low electron readout noise allow the upconversion scheme to surpass direct nuclear readout whenever $T_{\text{eff}}^{(e)}$ and $\tau_{a}$ are large enough to support coherent integration. Including both readout noise floors, the relative sensitivity is
$\frac{\text{SNR}_{e}}{\text{SNR}_{N}}=G_{\text{hyb}}\frac{\eta_{B}^{(N)}}{\eta_{B}^{(e)}}\sqrt{\frac{T_{\text{eff}}^{(e)}(\rho_{N})}{T_{\text{eff}}^{(N)}(\rho_{N})}}\,$, (4)
where the subscript eff indicates a nuclear density corrected effective coherence time derived from $T_{2}^{(e,N)}$, and $\eta^{(N,e)}_{B}$ are the magnetic sensitivity for the nuclear spin and electron spin respectively [1]. Thus hybrid electron readout with matched filtering exceeds direct nuclear spin detection when the hyperfine gain is appreciable, the electron readout noise is lower, and the electron coherence time is comparable to the nuclear coherence time.
Figure 1: The schematic shows a single unit cell; the experiment uses an ensemble/array of N coupled units. A gate-defined electron spin qubit (left) provides a fast, high-bandwidth readout channel, while a donor nuclear spin (right) serves as the axion-sensitive element. A static field $B_{0}=B_{z}\hat{z}$ defines the quantization axis. Gates P1, P2, P3, and P4 define the confinement potentials and are shown for illustration. Realizations can follow electron- and hole-based architectures in Refs. [47, 42, 31, 29, 36, 28, 30], hyperfine is local contact coupling at donor site with the donor nuclear spin acting as the probe of the axion wind field as in Refs. [23, 54, 55, 39, 12, 51].
Experimental proposal- The device shown in Fig. 1 is a chip-integrated hybrid nuclear electronic spin sensor designed for axion and ALP wind searches. All functional elements, including the donor ensemble, electrostatic control gates, microwave drive lines, and dispersive readout circuitry, are integrated on a single cryogenic platform. The chip operates inside radio-frequency and magnetic shielding to suppress ambient electromagnetic interference and low-frequency magnetic-field noise. Such noise sources are particularly critical because the axion signal is expected to exhibit sidereal modulation at frequency $\Omega_{\star}$ and annual sidebands at $\Omega_{\star}\pm\Omega_{\oplus}$. Suppressing environmental magnetic fluctuations at these frequencies is therefore essential to avoid systematic backgrounds that could mimic the expected modulation pattern. The device is thermally anchored to a dilution refrigerator stage at $T\lesssim 100~\mathrm{mK}$, which ensures long electron and nuclear coherence times, reduces thermal spin polarization noise, and stabilizes dispersive electron readout[22, 34]. Superconducting NbTi wiring together with filtered twisted-pair lines is used inside the shielded enclosure to suppress Johnson noise, attenuate high-frequency leakage, and minimize microwave phase noise in the control and readout chain. Under these conditions the dominant sensitivity limitation is set by quantum projection noise and amplifier noise rather than technical electromagnetic backgrounds.
While Si:209Bi maximizes the hyperfine-mediated gain due to its large $A=1.475$ GHz coupling, near-term experimental demonstrations of the upconversion principle could alternatively employ 28Si:31P donors. Although the smaller hyperfine interaction ($A=117$ MHz) reduces the hybrid gain in Eq. 3, which offers mature single donor control, high fidelity spin-to-charge readout, and well-established coherence benchmarks. A proof-of-principle validation of the hyperfine-mediated transduction and modulation fingerprint therefore does not require the maximal coupling strength of Bi and could be realized in existing 28Si:31P platforms before scaling to high-gain Bi ensembles.
Results Fig. 4 presents a forward-looking sensitivity projection that includes optional collective enhancement and resonator-assisted gain. The curves assume an ensemble size $N=10^{6}$ and a resonator gain $G_{\text{res}}$ corresponding to $Q=10^{5}$. These factors enter multiplicatively through $B_{e,\text{eff}}=F_{\text{ent}}G_{\text{res}}G_{\text{hyb}}B_{a}(N)$, so the discovery reach can be straightforwardly rescaled to other (near-term) values of $N$ and $Q$, and available quantum enhancement. Importantly, experimental validation of the hyperfine-mediated upconversion and modulation fingerprint does not require these enhancements and can be performed at the single-donor or small-array level. To estimate the axion-search sensitivity of the proposed sensors, we normalize the hybrid gain of each isotope to the sensitivity baseline introduced in Ref. [52]. Building on Eq. (3), we evaluate the electron phase filter $F_{e}(\omega)$ and the nuclear filter $Y_{N}(\omega)$ for a Ramsey/Hahn/CPMG sequence bank and determine the corresponding scan bands from their passbands. As expected from Eq. 4, the strong hyperfine coupling in 209Bi produces the largest hybrid gain at fixed $\tau_{N}$ and corresponding sensitivity advantage.
The hybrid spectral response can be written as $H(\omega;\tau,{N_{\pi}})=F_{e}(\omega;\tau)\,Y_{N}(\omega;\tau,{N_{\pi}})$, where $\tau$ is the sensing duration and $N_{\pi}$} the number of $\pi$ pulses. The electron filter $F_{e}$ suppresses low-frequency components and imposes fixed spectral zeros [Fig. 2(a)], while the nuclear filter $Y_{N}$ sets the tunable passband. Under CPMG control, the main passband is centered near $f_{c}\simeq{N_{\text{CPMG}}}/\tau$ with width of order $1/\tau$ and quality factor $Q\sim{N_{\text{CPMG}}}$ [Fig. 2(b)], enabling narrowband frequency targeting without sacrificing coherence. Together, $F_{e}$ and $Y_{N}$ define a controllable detection band that supports systematic scanning of the axion mass range.
Figure 2: (a) Normalized electron filter functions $|F(\omega)|^{2}/\tau^{2}$ for Ramsey and Hahn echo versus normalized frequency $\xi=f\tau$; the Hahn sequence narrows the main lobe and shifts the first zero from $\xi\simeq 1$ (Ramsey) to $\xi\simeq 2$. (b) Nuclear toggling filter amplitude $|Y_{N}(\omega)|/\tau$ for Ramsey, Hahn, and CPMG with ${N_{\text{CPMG}}}=4,8,16$, exhibiting passbands centered near $\xi\simeq{N_{\text{CPMG}}}$ with width $\sim 1/\tau$. This filter structure enables tunable frequency selectivity while preserving the sidereal modulation pattern, in contrast to conventional inductive NMR. Together these properties define the accessible mass-scan range.
Within this bounded window, matched filtering yields a resolved hybrid response with a signal-to-noise ratio scaling as $\mathrm{SNR}\propto A\sqrt{\tau}|H(\omega_{a})|$, where $H(\omega)$ denotes the joint electron–nuclear spectral filter. This structure enables mass scanning without degrading the sidereal modulation fingerprint of the axion wind. A full derivation of the scan constraints, pulse-selection rules, and the explicit bandwidth expressions used to generate Fig. 2 is provided in the Supplementary.
This scaling explains the monotonic trends also in Supplementary Material Fig. 1, where longer $T^{N}_{\text{eff}}$ directly lifts the hybrid gain, while larger hyperfine $A$, raises the transduction into the electron readout channel. Consistent with the modulation structure discussed earlier, the joint filter $H(\omega_{a})$ response preserves the same spectral fingerprint and therefore remains compatible with matched-filter detection.
Figure 3: (a) Time-domain evolution of the electron rotating-frame longitudinal magnetization $\langle S_{x}(\tau)\rangle$ under continuous spin lock at a Rabi frequency $\Omega_{R}/2\pi=50~\text{kHz}$. The axion-induced nuclear precession introduces a weak frequency modulation at $\omega_{a}/2\pi=5~\text{kHz}$. (b) Power spectral density of the same signal, $|\tilde{S}_{x}(\omega)|^{2}$, showing the locked carrier at $\Omega_{R}/2\pi$ (black dashed line) and symmetric axion-induced sidebands at $\Omega_{R}/2\pi\pm\omega_{a}/2\pi$ (red dashed lines). These spectral sidebands represent the frequency-encoded axion response characteristic of spin-lock detection.
While the pulsed protocols discussed above rely on accumulated phase to extract the axion-induced nuclear response, the same dynamics can also be accessed through a continuous-drive scheme. In this regime, the electronic qubit is held in a spin-locked state, so the axion-driven nuclear precession appears as a frequency modulation of the driven electron rather than a freely evolving phase. In the noiseless illustrative signal shown in Fig. 3, this manifests as weak modulation in the rotating-frame trajectory and as symmetric sidebands around the Rabi carrier, demonstrating the frequency-encoded detection mechanism without requiring a final phase-mapping pulse. This readout is not only a qualitative signature but also determines the architecture’s achievable sensitivity, since the spin-lock mode remains compatible with coherent integration and matched-filter analysis. The projected sensitivity enabled by this readout strategy is shown in Fig. 4.
Figure 4: Projected $5\sigma$ sensitivity to axion–nucleon couplings ($g_{an}=mg_{aNN}$). Calculated discovery thresholds on $g_{aNN}$ for a 209Bi donor ensemble using standard magnetometry protocols: Ramsey (Blue), Hahn-echo (Orange), XY8 (Green), and continuous spin-lock (Black). Curves include realistic device parameters and noise models [1], and assume collective enhancement from an entangled ensemble of $N=10^{6}$ nuclear spins plus resonator gain $G_{\mathrm{res}}(Q=10^{5})$. The effective signal field is $B_{e,\mathrm{eff}}=F_{\mathrm{ent}}G_{\mathrm{res}}G_{\mathrm{hyb}}B_{a}^{(N)}$, where $F_{\mathrm{ent}}$ parameterize quantum ensemble enhancement, e.g., $F_{\mathrm{ent}}=\sqrt{N}$ for uncorrelated spins at the SQL; $F_{\mathrm{ent}}=N$ for an ideal fully entangled state. Purple shading: Best spin-lock Monte-Carlo (95% C.L.) with the noise model in SM Tab.IV, giving best sensitivity over most masses. Yellow band: DFSZ model predictions; green Neutron star cooling region: current SN1987A cooling lower bounds. Portions of the parameter windows are from Ref. [46].
As determined by the filter-architecture constraints in Fig. 2, the achievable search region for the proposed hybrid platform is shown in Fig. 4. The Coloured band marks the portion of parameter space that is simultaneously compatible with the spectral passband and the hardware operating limits, representing the practically scannable mass window of the architecture. This window spans axion Compton frequencies from the sub-Hz regime to the kHz range, a range where conventional NMR-based wind searches are limited by inductive readout scaling. Regions outside this band are not eliminated physically, but lie outside the spectral response or operational range of the present implementation. The emergence of a contiguous, tunable, and readout-compatible scan band highlights a key advantage of the hybrid approach: it enables efficient access to the ultralight regime while maintaining signal contrast and spectral resolution, unlike purely nuclear or purely electronic sensing strategies.
Conclusion- In summary, we propose a hybrid nuclear–electronic upconversion scheme that exceeds the nuclear-inductive limit for axion-wind sensing while retaining directional (sidereal) signatures. Using a long-lived, axion-coupled nuclear ensemble and hyperfine transfer to a fast electronic readout, the method is compatible with silicon-donor solid-state qubit platforms and requires no new fundamental infrastructure. Monte Carlo simulations indicate the largest gain for bismuth donors, and with collective enhancement and resonator coupling we project $5\sigma$ DFSZ-level sensitivity over $m_{a}=10^{-16}-10^{-6}$ eV, reaching $g_{aNN}\sim 10^{-25}$.